(2xy^2 - 2y)dx + (3x^2y - 4x)dy = 0
1
Expert's answer
2020-11-26T19:54:42-0500
"Solution"

"\frac{∂}{∂y}(2xy^2 - 2y) - \frac{∂}{∂x}(3x^2y - 4x) = (4xy - 2) - (6xy - 4) = -2xy + 2."


Since "\frac{[\frac{∂}{∂y}(2xy^2 - 2y) - \frac{∂}{∂x}(3x^2y - 4x)] }{ (2xy^2 - 2y)}

=\frac{ -2(xy - 1)}{2y(xy - 1) }= \frac{-1}{y}" , a function of y alone,

we know that "e^{∫ (\frac{1}{y}) dy} = y" is an integrating factor.


Multiply both sides of the DE by "y" :

"(2xy^3 - 2y^2) dx + (3x^2y^2 - 4xy) dy = 0."


Now, this is an exact DE, because

"\frac{∂}{∂y}(2xy^3 - 2y^2) = 6xy^2 - 4y = \frac{∂}{∂x}(3x^2y^2 - 4xy)."


Since

"∫ (2xy^3 - 2y^2) dx = x^2 y^3 - 2xy^2 + f(y)\\

∫ (3x^2y^2 - 4xy) dy = x^2 y^3 - 2xy^2 + g(x),"

the general solution is "x^2 y^3 - 2xy^2 = C" .